I just need help setting this up.
A chemist has one solution that is 20% saline and a second that is 65% saline. The chemist makes 120 gallons of a solution that is 50% saline by combining the two solutions.
How much of each solution is used in the mixture?

Answer by stanbon(75887)

A chemist has one solution that is 20% saline and a second that is 65% saline. The chemist makes 120 gallons of a solution that is 50% saline by combining the two solutions.
---------------------------
Equation:
saline + saline = saline
0.20x + 0.65(120-x) = 0.50*120
Multiply thru by 100 to get:
20x + 65*120 - 65x = 50*120
-45x = -15*120
x = (1/3)120
x = 40 gallons (amt. of 20% solution needed in the mixture)
120-x = 120-40 = 80 gallons (amt. of 65% solution needed in the mixture)
====================================
Cheers,
Stan H.